# Grounded Question-Answering in Long Egocentric Videos

### v3

@@ Abstract
Existing approaches to video understanding, mainly designed for short videos from a third-person perspective, are limited in their applicability in certain fields, such as robotics. In this paper, we delve into open-ended question-answering (QA) in long, egocentric videos, which allows individuals or robots to inquire about their own past visual experiences. This task presents unique challenges, including the complexity of temporally grounding queries within extensive video content, the high resource demands for precise data annotation, and the inherent difficulty of evaluating open-ended answers due to their ambiguous nature. Our proposed approach tackles these challenges by (i) integrating query grounding and answering within a unified model to reduce error propagation; (ii) employing large language models for efficient and scalable data synthesis; and (iii) introducing a close-ended QA task for evaluation, to manage answer ambiguity. Extensive experiments demonstrate the effectiveness of our method, which also achieves state-of-the-art performance on the QAEgo4D and Ego4D-NLQ benchmarks. Code, data, and models are available at https://github.com/Becomebright/GroundVQA.
@@ 1 Introduction
Question-answering (QA) in long egocentric videos is challenging, primarily due to the complexity of temporally grounding and generate answers to the queries within extensive video content. A pioneer work, overlooking the importance of query grounding, achieves unsatisfactory QA performance that merely outperforms “blind guessing” [3]. On the other hand, research about temporal grounding in long egocentric videos, while achieving good progress, is limited in practical uses without the QA ability. A potential fix would be chaining models from the two areas, i.e., starting by localizing the temporal window to which the question relates, and followed by answering based on the corresponding video context. However, such method is often ineffective due to error propagation. To address these challenges, we propose to train a unified model for simultaneous query grounding and answering, as shown in Fig. 1. The unified training has three advantages: First, by training the grounding task, the model can better grasp query-relevant information from the long videos, which is helpful for effective QA; Second, simultaneously training these two tasks can reduce error accumulation thanks to the synergy effect [33] in deep models; Third, predicting a temporal window helps to understand the cause of failure while models generating wrong answers. Thus, we propose to solve the query grounding and answering concurrently, namely GroundVQA.
Nevertheless, training the unified architecture demands significant resources and effort to manually annotate the triplets – comprising a question, answer, and temporal window – on lengthy videos. Limited training data poses a major challenge in training large models with millions of parameters. To combat this issue, we establish an automatic pipeline that leverages large language models (LLMs) to generate abundant training samples. This pipeline prompts LLMs to transform the plentiful, timestamped narrations in Ego4D into QA pairs, and estimates corresponding temporal windows. As a result, we produce 303K data samples from 5,389 video clips, which is a 30-fold increase over the existing dataset [3]. Our newly created dataset, named EgoTimeQA, effectively mitigates overfitting and significantly enhances grounding and QA performance.
@@ 2 Related Work
Video language grounding. Video language grounding (VLG), initially proposed by Hendricks et al. [1], involves identifying and segmenting specific temporal intervals within third-person view videos based on a natural language description or query. Several datasets and benchmarks, such as Charades-STA[10] and TACoS [31], have been curated to support research in this field. Notably, the Ego4D-NLQ [13] dataset features long-form egocentric videos paired with natural language queries. The NaQ dataset [30] further expands NLQ by repurposing the extensive narrations within Ego4D as queries, thereby enhancing model performance [14]. However, these narrations are not directly applicable to question-answering (QA) tasks. To bridge this gap, we introduce an innovative generation pipeline that transforms these narrations into structured QA pairs. Additionally, we establish a multi-modal generative model capable of simultaneously localizing and localizing a language query given long, egocentric videos.
@@
Annotating VideoQA datasets is labor-intensive and expensive [40], while insufficient training data often results in over-fitting. To address this, automatic generation of VideoQA data has been investigated. For example, JustAsk [40] generates QA pairs from transcribed speech using a pre-trained language model for answer extraction and a separate model for question generation, substantially expanding the dataset size. More recently, Large language models (LLMs) have shown remarkable proficiency in task processing and reasoning. Innovative studies like LLaVA [23] and MiniGPT-4 [47] leverage the powerful capabilities of LLMs to generate visual instruction tuning data, achieving notable success in a range of visual-language tasks. In our study, we exploit LLMs to transform existing narrations from the Ego4D dataset into question-answer pairs with temporal windows, facilitating multimodal understanding for egocentric videos.
@@ 3.3 Generate QA from Narrations
Generating CloseQA data. We prompt the LLM to generate three options that appear valid but are ultimately incorrect for a given question-answer pair. The constructed prompt is illustrated in Fig. 3 (B). We apply this procedure to augment the EgoTimeQA dataset and the QaEgo4D benchmark, enabling the training and evaluation of models in a multi-choice scenario. The generation speed reaches 40K samples per hour.
Filtering CloseQA test set. The LLM may generate implausible choices. For example, some questions could be easily answered using dataset priors. To maintain the rigor of the CloseQA task, we filter out questions from the QaEgo4D test set that are easily answerable without video context. Specifically, we train a text-only “blind” model to identify and remove questions that are consistently answered correctly across ten trials with different seeds. The resulting QaEgo4D serves as a more refined testing ground. This ensures that models being evaluated truly require video content analysis to answer the questions correctly, thereby emphasizing the visual aspect of CloseQA.
@@ 4 Experiments
| Dataset | # Video | # Sample | Supported Task | |||
| OpenQA | CloseQA | VLG | ||||
train |
QaEgo4D | 997 | 11K | |||
| EgoTimeQA | 5,389 | 303K | ||||
@@ 4.1 Dataset and Metrics
EgoTimeQA is our contributed dataset, containing the same video clips as NaQ, while featuring 303K question-answer pairs with corresponding temporal windows.
QaEgo4D [3] expands the NLQ benchmark by manually annotating open-ended answers on its train and validation sets. It consists of 1,325 video clips and 14.5K data samples, further divided into 10,746 training, 1,913 validation, and 1,854 testing samples. It adopts Accuracy and machine translation metrics including ROUGE-L (f-score) [20], METEOR [2], and BLEU-4 [28]. In our experiments, we exclude BLEU-4 because the majority (around ) of answers in QaEgo4D are under three words in length, and Accuracy as it’s not effective due to language ambiguity in open-ended answers. To further address such ambiguity, we choose sentence similarity (Sim.) [32] as the primary metric, which maps sentences to a learned embedding space to calculate cosine similarity. Specifically, we utilize the Sentence Transformers library and the all-MiniLM-L6-v2 language model to perform the mapping.
@@ 4.3 QA Baselines
| Model | Additional Data | Additional Task | OpenQA | CloseQA | ||||
| EgoTimeQA | CloseQA | VLG | Sim. | ROUGE | METEOR | Accuracy | ||
| (A) | GroundVQA | – | – | – | 54.9 | 27.9 | 18.8 | – |
| (B) | GroundVQA | – | – | 54.8 | 27.7 | 18.7 | 36.50.7 | |
| (C) | GroundVQA | – | 55.6 | 29.0 | 19.8 | 37.50.2 | ||
| (D) | GroundVQA | – | 56.1 | 28.8 | 20.1 | 44.20.6 | ||
| (E) | GroundVQA | 57.7 | 30.2 | 21.2 | 45.61.2 | |||
@@
54.7 | 27.9 | 19.0 | 36.60.7 | |
| (J) | SimpleVQA+ |
55.4 | 28.1 | 19.5 | 39.50.8 | ||
| Training | EgoTimeQA | OpenQA | CloseQA | ||
| Sim. | ROUGE | METEOR | Accuracy | ||
| Two-stage | – | 54.7 | 27.3 | 18.4 | 36.60.6 |
| Unified | – | 55.6 | 29.0 | 19.8 | 37.50.2 |
@@
| Unified | 57.7 | 30.2 | 21.2 | 45.61.2 | |
| QaEgo4D | EgoTimeQA | NLQ+NaQ | Mean R@1 | Mean R@5 |
| – | – | 8.8 | 20.0 | |
| – | 18.4 | 37.2 | ||
| 20.9 | 42.5 |
@@ 4.4 Ablations
Integrating the CloseQA task. As presented in Tab. 2 (A-B), simultaneously training OpenQA and CloseQA tasks, despite their varying input-output formats, marginally impacts OpenQA performance. However, this integration offers a more comprehensive and reasonable method for assessing the model’s question-answering capabilities. Thus, we integrate CloseQA in training as default.
Integrating the VLG task. As shown in Tab. 2, incorporating VLG task indeed improves question-answering performance, e.g., GroundVQA’s Sentence Similarity increases from 54.8 to 55.6 when trained on QaEgo4D (B-C) and increases from 56.1 to 57.7 when trained on both QaEgo4D and EgoTimeQA (D-E), demonstrating the effectiveness of our proposed multi-task training approach.
@@
Unified v.s. separate training. An alternative to our unified model is training two separate models, one for temporal grounding and the other for question-answering on the grounded video clip. Results in Tab. 3 validate the effectiveness of our unified training method.
Incorporating EgoTimeQA data. Our data generation method produces 303K samples, a 30-fold increase over the QaEgo4D training set, resulting in notable performance gains in QA and VLG tasks. The QA metrics for GroundVQA demonstrate significant enhancements, as evidenced in Tab. 2 (B to D and C to E). Similar improvements are observed for SimpleVQA (G-H) and SimpleVQA+ (I-J), confirming the generality and effectiveness for EgoTimeQA. In Tab. 4, EgoTimeQA also boosts VLG recall by a large margin, which is further amplified with NLQ and NaQ data. Notably, as depicted in Fig. 4, the value of EgoTimeQA is even more evident in overcoming overfitting.
Combining the above enhancements (A-E in Tab. 2), our method closely approaches the oracle upper bound (F), with the main gap due to imperfect temporal grounding.
@@ 4.5 Comparison with State-of-the-art
On QaEgo4D. We report results on the QaEgo4D test set in Tab. 5. To ensure fairness, we reproduce the other methods using identical settings (detailed in Sec. 4.3). Our model achieves the best performance, outperforming prior works by a large margin.
On NLQ. We then assess VLG performance on the NLQ test set. As seen in Tab. 6, though it’s not our primary focus, our model, without complex design or multi-scale feature pyramids [14], matches state-of-the-art performance. GroundVQA exhibits further improvements by pre-training on NLQ and NaQ, and fine-tuning exclusively on NLQ.
| Method | OpenQA | CloseQA | Param | ||
| Sim. | ROUGE | METEOR | Accuracy | ||
@@
| RM | 54.1 | 27.3 | 18.5 | 38.10.9 | 368 |
| GroundVQA | 58.2 | 30.4 | 21.5 | 47.60.8 | 252 |
Qualitative analysis. Fig. 5(A) shows an OpenQA example. Our model successfully predicts the temporal window and the answer, while SimpleVQA fails. Although our predicted answer slightly differs from the ground truth, it’s still valid, highlighting the challenge of paraphrasing in evaluating open-ended answers, thus reflecting the advantage of our CloseQA task. Fig. 5(B) demonstrates a CloseQA example. Our model shows competence in predicting a close temporal window and identifying the correct answer. On the contrary, SimpleVQA chooses an incorrect answer, while the absence of temporal localization hinders understanding of its error source. Fig. 5(C) is a failure case of our model and SimpleVQA. Yet, our model’s temporal window prediction is relevant to the query, and the predicted answer is coherent with the grounded content. This highlights an issue of the QaEgo4D and NLQ annotations, where multiple relevant video segments and plausible answers exist, but only one annotation is available per query.
| Method | Recall@1 | Recall@5 | |||
| Mean | IoU=0.3 | IoU=0.5 | IoU=0.3 | IoU=0.5 | |
| VSLNet [44] | 4.08 | 5.42 | 2.75 | 8.79 | 5.07 |
| EgoVLP [21] | 8.35 | 10.46 | 6.24 | 16.76 | 11.29 |
| ReLER [24] | 10.51 | 12.89 | 8.14 | 15.41 | 9.94 |
| NaQ++ [30] | 17.67 | 21.70 | 13.64 | 25.12 | 16.33 |
| GroundNLQ [14] | 20.91 | 24.50 | 17.31 | 40.46 | 29.17 |
| GroundVQA | 19.31 | 23.65 | 14.96 | 36.19 | 24.58 |
@@
We present additional experimental results in the supplementary material, including the impact of using different LLMs to generate QA data, a more in-depth statistical analysis of EgoTimeQA, further qualitative findings, etc.
@@ 5 Conclusion
Appendix
@@
In the main paper, we default to using Llama2-13B-chat for generating QA data. Here, we experiment with ChatGPT-3.5-turbo111Utilizing the OpenAI API: https://platform.openai.com. To expedite the data generation and model training process, we reduce the amount of data relative to EgoTimeQA. Specifically, we use both LLMs to generate QA data from the NLQ training set, which includes 1.3K video clips and 221K narration sentences. Full prompts are detailed in Sec. D, where minor differences exist between ChatGPT’s and Llams2’s prompts. Consequently, Llama2 produces 92K and ChatGPT 97K data pairs.222The variation in numbers stems from differing rates of generation errors, i.e., the generated string cannot be converted into a dictionary containing “Q” and “A” as the in-context examples. Compared to QaEgo4D, the video clips are almost identical, but the QA pairs are denser in time. We then train GroundVQA on each dataset to assess generation quality, referencing OpenQA and VLG performance. Note that for CloseQA, both training and testing data are generated by the LLMs. Thus, evaluating CloseQA on a test set produced by one LLM, such as Llama-2, would be unfair for ChatGPT because of the bias in the generation process. Therefore, we exclude CloseQA evaluation from this experiment.
@@
| Training Data | OpenQA | VLG | |||||||
| Source | # Clip | # Sample | Cost | Sim. | ROUGE | METEOR | Mean R@1 | Mean R@5 | |
| (A) | QaEgo4D | 1.0K | 11K | – | 55.6 | 29.0 | 19.8 | 8.8 | 20.0 |
| (B) | + ChatGPT QA (NLQ) | 1.3K | 107K | $50 | 56.9 | 29.2 | 19.8 | 15.7 | 33.7 |
| (C) | + Llama2 QA (NLQ) | 1.3K | 103K | 5 Gh | 57.1 | 29.4 | 20.3 | 16.3 | 34.3 |

### v4

@@ Abstract
Existing approaches to video understanding, mainly designed for short videos from a third-person perspective, are limited in their applicability in certain fields, such as robotics. In this paper, we delve into open-ended question-answering (QA) in long, egocentric videos, which allows individuals or robots to inquire about their own past visual experiences. This task presents unique challenges, including the complexity of temporally grounding queries within extensive video content, the high resource demands for precise data annotation, and the inherent difficulty of evaluating open-ended answers due to their ambiguous nature. Our proposed approach tackles these challenges by (i) integrating query grounding and answering within a unified model to reduce error propagation; (ii) employing large language models for efficient and scalable data synthesis; and (iii) introducing a close-ended QA task for evaluation, to manage answer ambiguity. Extensive experiments demonstrate the effectiveness of our method, which also achieves state-of-the-art performance on the QAEgo4D and Ego4D-NLQ benchmarks. Code, data, and models are open-sourced 111https://github.com/Becomebright/GroundVQA.
@@ 1 Introduction
Question-answering (QA) in long egocentric videos is challenging, primarily due to the complexity of temporally grounding and generating answers to the queries within extensive video content. A pioneer work, overlooking the importance of query grounding, achieves unsatisfactory QA performance that merely outperforms “blind guessing” [3]. On the other hand, research about temporal grounding in long egocentric videos, while achieving good progress, is limited in practical uses without the QA ability. A potential fix would be chaining models from the two areas, i.e., starting by localizing the temporal window to which the question relates, and followed by answering based on the corresponding video context. However, such a method is often ineffective due to error propagation. To address these challenges, we propose to train a unified model for simultaneous query grounding and answering, as shown in Fig. 1. The unified training has three advantages: First, by training the grounding task, the model can better grasp query-relevant information from the long videos, which is helpful for effective QA; Second, simultaneously training these two tasks can reduce error accumulation thanks to the synergy effect [35] in deep models; Third, predicting a temporal window helps to understand the cause of failure. Thus, we propose to solve the query grounding and answering concurrently, namely GroundVQA.
Nevertheless, training the unified architecture demands significant resources and effort to manually annotate the triplets – comprising a question, answer, and temporal window – on lengthy videos. Limited training data poses a major challenge in training large models with millions of parameters. To combat this issue, we establish an automatic pipeline that leverages large language models (LLMs) to generate abundant training samples. This pipeline prompts LLMs to transform the plentiful, timestamped narrations in Ego4D into QA pairs, and estimates corresponding temporal windows. As a result, we produce 303K data samples from 5,389 video clips, which is a 30-fold increase over the existing dataset [3]. Our newly created pre-training dataset, named EgoTimeQA, effectively mitigates overfitting and significantly enhances grounding and QA performance.
@@ 2 Related Work
Video language grounding. Video language grounding (VLG), initially proposed by Hendricks et al. [1], involves identifying and segmenting specific temporal intervals within third-person view videos based on a natural language description or query [48, 42, 22]. Several datasets and benchmarks, such as Charades-STA[10] and TACoS [33], have been curated to support research in this field. Notably, the Ego4D-NLQ [13] dataset features long-form egocentric videos paired with natural language queries. The NaQ dataset [32] further expands NLQ by repurposing the extensive narrations within Ego4D as queries, thereby enhancing model performance [14]. However, these narrations are not directly applicable to question-answering (QA) tasks. To bridge this gap, we introduce a generation pipeline that transforms these narrations into structured QA pairs. Additionally, we establish a multi-modal generative model capable of temporally localizing and answering a language query given long, egocentric videos.
@@
Annotating VideoQA datasets is labor-intensive and expensive [43], while insufficient training data often results in over-fitting. To address this, automatic generation of VideoQA data has been investigated. For example, JustAsk [43] generates QA pairs from transcribed speech using pre-trained language models, substantially expanding the dataset size. More recently, Large language models (LLMs) have shown remarkable proficiency in task processing and reasoning. Innovative studies like LLaVA [24] and MiniGPT-4 [51] leverage the powerful capabilities of LLMs to generate visual instruction tuning data, achieving notable success in a range of visual-language tasks. In our study, we exploit LLMs to transform existing narrations from the Ego4D dataset into question-answer pairs with temporal windows, facilitating multimodal understanding for egocentric videos. A concurrent work, EgoSchema [28], also exploits LLMs for constructing QA pairs. Compared to it, our approach includes both CloseQA and OpenQA, offering greater real-world applicability. Moreover, EgoSchema aims to summarize entire videos, while our method emphasizes episodic memory, focusing on recalling specific fragments for fine-grained queries.
@@ 3.3 Generate QA from Narrations
Generating CloseQA data. We prompt the LLM to generate three options that appear valid but are ultimately incorrect for a given question-answer pair. The constructed prompt is illustrated in Fig. 3 (B). We apply this procedure to augment EgoTimeQA and QaEgo4D, enabling the training and evaluation of models in a multi-choice scenario. The generation speed reaches 40K samples per hour.
Filtering CloseQA test set. The LLM may generate implausible choices. To maintain the rigor of the CloseQA task, we filter out questions from the QaEgo4D test set that are easily answerable without video context. Specifically, we train a text-only “blind” model to identify and remove questions that are consistently answered correctly across ten trials with different seeds. Additionally, we perform rigorous human verification by eliminating samples that contain incorrect answers or temporal windows. The resulting QaEgo4D serves as a more refined testing ground. This ensures that models being evaluated truly require video content analysis to answer the questions correctly, thereby emphasizing the visual aspect of CloseQA.
@@ 4 Experiments
Dataset | # Video | # Sample | Supported Task | ||||
OpenQA | CloseQA | VLG | |||||
train | QaEgo4D | 997 | 11K | ||||
EgoTimeQA | 5,389 | 303K | |||||
@@ 4.1 Dataset and Metrics
EgoTimeQA is our contributed pre-training dataset, containing the same video clips as NaQ, while featuring 303K question-answer pairs with temporal windows.
QaEgo4D [3] expands the NLQ benchmark by manually annotating open-ended answers on its train and validation sets. It consists of 1,325 video clips and 14.5K data samples, further divided into 10,746 training, 1,913 validation, and 1,850 testing samples. It adopts Accuracy and machine translation metrics including ROUGE-L (f-score) [20], METEOR [2], and BLEU-4 [30]. In our experiments, we exclude BLEU-4 because the majority (around ) of answers in QaEgo4D are under three words in length, and Accuracy as it’s not effective due to language ambiguity in open-ended answers. To further address such ambiguity, we choose sentence similarity (Sim.) [34] as the primary metric, which maps sentences to a learned embedding space to calculate cosine similarity. Specifically, we utilize the Sentence Transformers library and the all-MiniLM-L6-v2 language model to perform the mapping.
@@ 4.3 QA Baselines
cl c cccc ccc cc
Model Additional Data Additional Task OpenQA CloseQA
EgoTimeQA CloseQA VLG Sim. ROUGE METEOR Accuracy
(A) GroundVQA – – – 54.9 27.9 18.8 –
(B) GroundVQA – – 54.8 27.7 18.7 39.50.5
(C) GroundVQA – 55.6 29.0 19.8 40.81.0
(D) GroundVQA – 56.1 28.8 20.1 47.20.5
(E) GroundVQA 57.7 30.2 21.2 48.70.4
@@
54.7 27.9 19.0 39.30.6
(J) SimpleVQA+
55.4 28.1 19.5 42.00.7
lc ccccc c
Training EgoTimeQA OpenQA CloseQA
Sim. ROUGE METEOR Accuracy
Two-stage – 54.7 27.3 18.4 39.30.8
Unified – 55.6 29.0 19.8 40.81.0
@@
Unified 57.7 30.2 21.2 48.70.4
ccc—cc
QaEgo4D EgoTimeQA NLQ+NaQ Mean R@1 Mean R@5
– – 8.8 20.0
– 18.4 37.2
20.9 42.5
@@ 4.4 Ablations
Integrating the CloseQA task. As presented in Tab. 4.3 (A-B), simultaneously training OpenQA and CloseQA tasks, despite their varying input-output formats, marginally impacts OpenQA performance. However, this integration offers a more comprehensive and reasonable method for assessing the model’s question-answering capabilities. Thus, we integrate CloseQA in training as default.
Integrating the VLG task. As shown in Tab. 4.3, incorporating VLG task indeed improves question-answering performance, e.g., GroundVQA’s Sentence Similarity increases from 54.8 to 55.6 when trained on QaEgo4D (B-C) and increases from 56.1 to 57.7 when trained on both QaEgo4D and EgoTimeQA (D-E), demonstrating the effectiveness of our proposed multi-task training approach.
@@
Unified v.s. separate training. An alternative to our unified model is training two separate models, one for temporal grounding and the other for question-answering on the grounded video clip. Results in Tab. 4.3 validate the effectiveness of our unified training method.
Incorporating EgoTimeQA data. Our data generation method produces 303K samples, a 30-fold increase over the QaEgo4D training set, resulting in notable performance gains in QA and VLG tasks. The QA metrics for GroundVQA demonstrate significant enhancements, as evidenced in Tab. 4.3 (B to D and C to E). Similar improvements are observed for SimpleVQAS (G-H) and SimpleVQA+S (I-J), confirming the generality and effectiveness for EgoTimeQA. In Tab. 4.3, EgoTimeQA also boosts VLG recall by a large margin, which is further amplified with NLQ and NaQ data. Notably, as depicted in Fig. 4, the value of EgoTimeQA is even more evident in overcoming overfitting.
Combining the above enhancements (A-E in Tab. 4.3), our method closely approaches the oracle upper bound (F), with the main gap due to imperfect temporal grounding.
@@ 4.5 Comparison with State-of-the-art
On QaEgo4D. We report results on the QaEgo4D test set in Tab. 4.5. To ensure fairness, we reproduce the other methods using identical settings (detailed in Sec. 4.3). Our model achieves the best performance, outperforming prior works by a large margin.
On NLQ. We then assess VLG performance on the NLQ test set. As seen in Tab. 6, our model, without complex design or multi-scale feature pyramids [14], matches the SOTA performance. GroundVQA exhibits further improvements by pre-training on NLQ and NaQ, and fine-tuning exclusively on NLQ.
l ccc c c cc
Method OpenQA CloseQA Param
Sim. ROUGE METEOR Accuracy
@@
RM∗ 54.1 27.3 18.5 39.90.8 368
GroundVQA 58.2 30.4 21.5 50.20.5 252
Qualitative analysis. Fig. 5(A) shows an OpenQA example. Our model successfully predicts the temporal window and the answer, while SimpleVQA∗ fails. Although our predicted answer slightly differs from the ground truth, it’s still valid, highlighting the challenge of paraphrasing in evaluating open-ended answers, thus reflecting the advantage of our CloseQA task. Fig. 5(B) demonstrates a CloseQA example. Our model shows competence in predicting a close temporal window and identifying the correct answer. On the contrary, SimpleVQA∗ chooses an incorrect answer, while the absence of temporal localization hinders understanding of its error source. Fig. 5(C) is a failure case of our model and SimpleVQA∗. Yet, our model’s temporal window prediction is relevant to the query, and the predicted answer is coherent with the grounded content. This highlights an issue of the QaEgo4D and NLQ annotations, where multiple relevant video segments and plausible answers exist, but only one annotation is available per query.
Method | Recall@1 | Recall@5 | |||||
Mean | IoU=0.3 | IoU=0.5 | IoU=0.3 | IoU=0.5 | |||
VSLNet [47] | 4.08 | 5.42 | 2.75 | 8.79 | 5.07 | ||
EgoVLP [21] | 8.35 | 10.46 | 6.24 | 16.76 | 11.29 | ||
ReLER [25] | 10.51 | 12.89 | 8.14 | 15.41 | 9.94 | ||
NaQ++ [32] | 17.67 | 21.70 | 13.64 | 25.12 | 16.33 | ||
GroundNLQ [14] | 20.91 | 24.50 | 17.31 | 40.46 | 29.17 | ||
GroundVQA | 19.31 | 23.65 | 14.96 | 36.19 | 24.58 | ||
@@
We present additional results in the supplementary material, including the impact of using different LLMs to generate QA data, a more in-depth statistical analysis of EgoTimeQA, additional qualitative findings, prompts for generating QA data, limitations, and future work.
@@ 5 Conclusion
Grounded Question-Answering in Long Egocentric Videos
Supplementary Material
@@
In the main paper, we default to using Llama2-13B-chat for generating QA data. Here, we experiment with ChatGPT-3.5-turbo222Utilizing the OpenAI API: https://platform.openai.com. To expedite the data generation and model training process, we reduce the amount of data relative to EgoTimeQA. Specifically, we use both LLMs to generate QA data from the NLQ training set, which includes 1.3K video clips and 221K narration sentences. Full prompts are detailed in Sec. D, where minor differences exist between ChatGPT’s and Llams2’s prompts. Consequently, Llama2 produces 92K and ChatGPT 97K data pairs.333The variation in numbers stems from differing rates of generation errors, i.e., the generated string cannot be converted into a dictionary containing “Q” and “A” as the in-context examples. Compared to QaEgo4D, the video clips are almost identical, but the QA pairs are denser in time. We then train GroundVQA on each dataset to assess generation quality, referencing OpenQA and VLG performance. Note that for CloseQA, both training and testing data are generated by the LLMs. Thus, evaluating CloseQA on a test set produced by one LLM, such as Llama-2, would be unfair for ChatGPT because of the bias in the generation process. Therefore, we exclude CloseQA evaluation from this experiment.
@@
Training Data | OpenQA | VLG | |||||||||
Source | # Clip | # Sample | Cost | Sim. | ROUGE | METEOR | Mean R@1 | Mean R@5 | |||
(A) | QaEgo4D | 1.0K | 11K | – | 55.6 | 29.0 | 19.8 | 8.8 | 20.0 | ||
(B) | + ChatGPT QA (NLQ) | 1.3K | 107K | $50 | 56.9 | 29.2 | 19.8 | 15.7 | 33.7 | ||
(C) | + Llama2 QA (NLQ) | 1.3K | 103K | 5 Gh | 57.1 | 29.4 | 20.3 | 16.3 | 34.3 | ||